OpenStudy (anonymous):

The period of f(x)=cos (2pi/3)x is

OpenStudy (anonymous):

\[\cos (2\pi/3)x\]

OpenStudy (anonymous):

cos x has a period of 2pi. with a multiple of K this will have a multiple of 1/k. so, 3/2pi . 2pi ie 3

OpenStudy (anonymous):

if you're specific about the rigor, period means that \[\cos ( 2\Pi \div3 (x+\alpha)) = \cos ( 2\Pi \div3 (x))\] . this gives \[2 \pi \div 3 \alpha = 2 \pi\] . so \[\alpha\] equals 3. Cheers!

OpenStudy (anonymous):

thanks!!